How many feet will a car travel in one second at 6 mph?

Answer by Maths68(1474)

How many feet will a car travel in one second at 6 mph?
Car speed = 6 mph ..........(1)
Change miles into feet and hour into second
1 mile = 5082 feet
Multiply by 6 both sides of above equation
6*1 mile=6*5082 feet
6 miles = 30492 feet

1 hour = 1*60*60 = 3600 seconds
Put above values of 1 hour and 6 miles in (1)
Car speed = 6 mph ..........(1)
Car speed = 30492/3600
Car speed = 8.47 feet/second